第 13 回福岡数論研究集会

本研究集会は, 一部日本学術振興会科学研究費基盤研究 (S) 16H06336, 挑戦的萌芽研究 18K18712 （研究代表者：金子昌信）の補助を受けています。

日時：2019 年 8 月 7 日 (水) − 8 月 9 日 (金)

場所：九州大学（伊都キャンパス）講義棟３０２（ビックリーフ 3 階）福岡市西区元岡 744

プログラム

8 月 7 日 (水)

9 : 55 − 10 : 00 Opening
10 : 00 − 10 : 45 齊藤暢（中央大学）

幾何的視点から観た Lucas 数列〜3 階の場合

11 : 00 − 11 : 45 藤井大輔（名古屋大学）

一般化 3x + 1 問題の軌道周期

13 : 15 − 14 : 00 岡崎勝男（九州大学）

高さ 1 の行列

14 : 15 − 15 : 00 渋川元樹（神戸大学）

Multivariate Bernoulli polynomials

15 : 15 − 16 : 00 村上和明（慶應義塾女子高等学校）

On the cyclicity of the unramified Iwasawa modules of the maximal multiple ℤ_{p}-extensions over imaginary quadratic fields

16 : 15 − 17 : 00 植木潤（東京電機大学）

捻じれ Alexander 多項式の副有限剛性について

8 月 8 日 (木)

10 : 00 − 10 : 45 村上拓也（九州大学）

On Hoffman’s t-values of maximal height

11 : 00 − 11 : 45 小野雅隆（九州大学）

シャッフル型対称多重ゼータ値の級数表示

13 : 15 − 14 : 00 中屋智瑛（九州大学）

古典的楕円関数と保型微分方程式について

14 : 15 − 15 : 00 小谷久寿（Max Planck Institute for Mathematics）

On certain multiple zeta functions for a flat bundle over a Riemannian manifold

15 : 15 − 16 : 00 源嶋孝太（京都産業大学）

ランキン・セルバーグ級数の特殊値に関するある平均について

16 : 15 − 17 : 00 軍司圭一（千葉工業大学）

On the Fourier coefficients of Siegel Eisenstein series and genus theta series

18 : 30 −  懇親会

8 月 9 日 (金)

10 : 00 − 10 : 45 石本宙（京都大学）

Local intertwining relation for metaplectic groups

11 : 00 − 11 : 45 奥村喜晶（東京工業大学）

Drinfeld 加群の Q-仮想性について

13 : 15 − 14 : 00 長町一平（東京大学）

On the Hom version of the Grothendieck Conjecture for morphisms from regular varieties to hyperbolic polycurves of dimension 2

14 : 15 − 15 : 00 佐藤信夫（九州大学）

Deligne-Glanois type basis for the alternating multiple zeta values

15 : 15 − 16 : 00 広瀬稔（九州大学）

Multivariable generalizations of Zhao’s generalized 2-1 formula and Zagier’s 2-3-2 formula

16 : 00 − 16 : 05 Closing

※ 研究集会の最新情報は下記アドレスでご確認ください。

  http://auemath.aichi-edu.ac.jp/~ykishi/FSNT/19-fukuokaNT.html

世話人：金子昌信（九州大学) mkaneko[at]math.kyushu-u.ac.jp
権寧魯（九州大学) ygon[at]math.kyushu-u.ac.jp
岸康弘（愛知教育大学)ykishi[at]auecc.aichi-edu.ac.jp